# When in Doubt, Deliberate: Confidence-Based Routing to Expert Debate for Sexism Detection

### v1

@@ A.2.2 Binary Classification Optimizations
Class-balanced cross-entropy (CB-CE) is emplyed to address class imbalance by re-weighting loss contributions based on effective sample counts. For each class $y$ with $n_{y}$ training examples, the effective number accounts for information overlap:
@@ Appendix B Prompts
This section presents the prompt templates used throughout In the CEJ framework. Initial persona judgments are collected using the prompt shown in Figure A4. This is followed by the structured debate prompt (Figure A5), which enables each persona to reflect on peer opinions, revise its stance if persuaded, and recalibrate its confidence. The discussion is then summarized and passed to the judge prompt (Figure A6), which concludes the process by generating a final label along with a justification and calibrated confidence score. All prompts rely on consistent definitions and structured output formats (e.g., JSON objects with fields such as label, justification, and confidence) to support interpretability and downstream evaluation.

### v2

@@ A.2.2 Binary Classification Optimizations
Class-balanced cross-entropy (CB-CE) is employed to address class imbalance by re-weighting loss contributions based on effective sample counts. For each class $y$ with $n_{y}$ training examples, the effective number accounts for information overlap:
@@ Appendix B Prompts
This section presents the prompt templates used throughout the CEJ framework. Initial persona judgments are collected using the prompt shown in Figure A4. This is followed by the structured debate prompt (Figure A5), which enables each persona to reflect on peer opinions, revise its stance if persuaded, and recalibrate its confidence. The discussion is then summarized and passed to the judge prompt (Figure A6), which concludes the process by generating a final label along with a justification and calibrated confidence score. All prompts rely on consistent definitions and structured output formats (e.g., JSON objects with fields such as label, justification, and confidence) to support interpretability and downstream evaluation.